### v1

# Trust in Autonomous Human–Robot Collaboration: Effects of Responsive Interaction Policies

Abstract
Trust plays a central role in human–robot collaboration, yet its formation is rarely examined under the constraints of fully autonomous interaction. This pilot study investigated how interaction policy influences trust during in-person collaboration with a social robot operating without Wizard-of-Oz control or scripted repair. Participants completed a multi-stage collaborative task with a mobile robot that autonomously managed spoken-language dialogue, affect inference, and task progression. Two interaction policies were compared: a responsive policy, in which the robot proactively adapted its dialogue and assistance based on inferred interaction state, and a neutral, reactive policy, in which the robot provided only direct, task-relevant responses when prompted. Responsive interaction was associated with significantly higher post-interaction trust under viable communication conditions, despite no reliable differences in overall task accuracy. Sensitivity analyses indicated that affective and experiential components of trust were more sensitive to communication breakdown than evaluative judgments of reliability, and that as language-mediated interaction degraded, the trust advantage associated with responsiveness attenuated and ratings became less clearly interpretable as calibrated evaluations of collaborative competence. These findings suggest that trust in autonomous human–robot interaction emerges from process-level interaction dynamics and operates within constraints imposed by communication viability, highlighting the importance of evaluating trust under real autonomy conditions when designing interactive robotic systems.
Keywords human-robot collaboration • HRI • socially assistive robotics • cobots • autonomous robot systems • embodied AI • spoken language interaction • trust in human-robot interaction • affect-adaptive robot systems
1 Introduction
As artificial intelligence (AI) technologies advance, they are increasingly integrated into robotic systems, enabling more adaptive, autonomous, and context-sensitive behaviour in real-world environments. This convergence has accelerated the deployment of robots across safety-critical domains such as manufacturing [11], mining [18], and healthcare [12], where robots are now expected to operate alongside humans rather than in isolation [15, 37]. In these collaborative settings, successful deployment depends not only on technical performance and safety guarantees, but also on whether human users are willing to rely on, communicate with, and coordinate their actions around robotic partners [8, 16]. Trust has therefore emerged as a central determinant of adoption and effective use in human–robot collaboration (HRC) [41, 8]. Insufficient trust can lead to disuse or rejection of automation, while excessive trust risks overreliance and accidents—particularly in environments characterized by uncertainty or incomplete information [13].
In HRC, trust is commonly understood as a willingness to rely on an agent under conditions of uncertainty and risk [30, 21]. This reliance is dynamically calibrated, shaping how closely users monitor a robot, when they intervene, and whether they defer or override its actions. Appropriately calibrated trust supports effective coordination, whereas under-trust may result in disengagement or redundant oversight, and over-trust can lead to inappropriate reliance and unsafe outcomes [13]. These dynamics are especially pronounced in dialogue-driven collaborative tasks, where misunderstandings, delays, or ambiguous responses may directly influence users’ ongoing assessments of a robot’s competence and reliability.
A substantial body of human–robot interaction (HRI) research has examined how robot behaviour influences user trust, perceived reliability, and cooperation across industrial and social contexts [35, 17]. Trust is typically conceptualized as a multidimensional construct encompassing cognitive evaluations of competence, predictability, and reliability, alongside a behavioural willingness to collaborate toward shared goals under conditions of risk or uncertainty [30, 21, 13]. Despite this multidimensional framing, empirical studies have most often operationalized trust using post-interaction self-report questionnaires collected following short, highly controlled, and often scripted interactions. While such measures provide valuable global assessments of user attitudes, they offer limited insight into how trust is negotiated, disrupted, and repaired during ongoing interaction—particularly in autonomous systems where errors and ambiguities are unavoidable [26].
Studying trust as an interactional process therefore requires experimental settings in which users engage with robots that exhibit both adaptive behaviour and realistic system limitations [8]. In such settings, trust is shaped not only by task success but by how robots handle uncertainty, errors, and misalignment during interaction. Fully autonomous systems, where dialogue management and response generation occur without human intervention, provide a critical testbed for examining these dynamics, as they expose users to the same constraints and breakdowns encountered in real-world deployment [8].
Yet in practice, much of the existing HRI trust literature has relied on scripted behaviours, simulated environments, or Wizard-of-Oz paradigms in which a human operator covertly manages the robot’s behaviour [4, 8]. While these approaches are valuable for isolating specific design factors, they obscure the interaction breakdowns and system imperfections that characterize deployed autonomous robots. Limitations such as speech recognition errors, delayed responses, misinterpretations of user intent, and incomplete affect sensing are not peripheral issues but defining features of real-world interaction. These failures are likely to play a decisive role in shaping trust and collaboration, yet remain underrepresented in empirical evaluations [8].
Within HRI, a range of design strategies have been proposed to support appropriate trust calibration during collaboration, including robot appearance [31], transparency cues [42], explanations of system intent, adaptive feedback [38], and interaction pacing [22]. Many of these approaches aim to help users form accurate expectations about a robot’s capabilities and limitations, particularly in contexts involving uncertainty or partial observability. Among these strategies, adaptive interaction behaviour—how and when a robot responds to user state and task context—has been identified as a particularly influential factor in shaping perceptions of competence, reliability, and collaboration [17].
Recent advances in AI have expanded the range of interaction strategies available to autonomous robotic platforms in practice, enabling systems to move beyond fixed, scripted behaviours toward adaptive interaction policies that respond to user state and task context in real time [2]. Improvements in spoken-language processing, dialogue state tracking, and large language model–based reasoning now allow robots to adjust not only what they say, but when and how assistance is provided during collaboration [39]. In parallel, advances in affect inference from language and interaction cues have made it increasingly feasible for robots to incorporate estimates of user emotional state into interaction management [27, 37]. As a result, responsiveness in contemporary HRI is increasingly understood as a property of an underlying interaction policy, governing how a robot interprets cues, initiates support, and manages uncertainty, rather than as a surface-level social behaviour [5, 35, 17].
From an engineering perspective, responsiveness in autonomous human–robot interaction is not implemented as a single behavioural rule or surface-level cue, but emerges from the coordination of multiple system components responsible for perception, interpretation, and action [1]. In practice, responsive behaviour depends on the integration of dialogue management, state tracking, and inference mechanisms that estimate task progress, interaction uncertainty, and user affect. Proactive assistance based on this integrated state representation differs fundamentally from reactive, request-based behaviour; for example, a robot may offer clarification, encouragement, or pacing adjustments when confusion, hesitation, or frustration is inferred, rather than waiting for an explicit request for help [5].
Implementing such behaviour requires autonomous systems to manage spoken-language dialogue, maintain interaction state over time, and coordinate verbal and nonverbal responses in real time, all while operating under noise, latency, and sensing uncertainty [8]. As a result, responsiveness in deployed systems reflects properties of the overall interaction architecture, not merely the presence or absence of adaptive dialogue. Despite growing interest in responsive and affect-aware robots, relatively little empirical work has examined how such integrated interaction systems operate under fully autonomous conditions, or how their behaviour shapes trust and collaboration when communication breakdowns cannot be externally repaired [8].
The present pilot study addresses these gaps by examining trust and collaboration during in-person interaction with a fully autonomous social robot in which participants collaborated with one of two versions of the same robot platform during a dialogue-driven puzzle task requiring shared problem solving. In both conditions, all interaction management—including speech recognition, dialogue state tracking, task progression, and response generation was handled and logged autonomously by a centralized dialogue agent without human intervention. In the responsive condition, the robot employed a proactive, affect-aware interaction policy, adapting its assistance based on conversational cues and inferred user affect (e.g., frustration or engagement). In the neutral condition, the robot followed a reactive policy, providing basic guidance and periodic check-ins, otherwise providing assistance only when explicitly requested.
This pilot study had three primary objectives: (1) to design and evaluate the feasibility of an autonomous spoken-language interaction system with affect-responsive behaviour on a mobile robot platform; (2) to assess whether a responsive interaction policy influences post-interaction trust and collaborative experience under autonomous conditions; and (3) to explore how behavioural and interaction-level indicators align with subjective trust evaluations.
Rather than optimizing for flawless interaction, this pilot study prioritizes feasibility assessment and mechanism exploration over confirmatory hypothesis testing. The sample size is modest, interaction conditions subject to real-world autonomy constraints, and final condition assignment was not fully randomized due to participant attrition. Accordingly, results should be interpreted as exploratory evidence regarding how interaction policy shapes trust under autonomous spoken-language collaboration, and as guidance for the design of larger, confirmatory studies. The findings are intended to inform the design of a larger subsequent study by evaluating feasibility and identifying technical, interactional, and methodological challenges that must be addressed when evaluating affect-responsive robots in real-world contexts.
2 Methods
This study employed a between-subjects experimental design to examine how robot interaction policy influences trust and collaboration during fully autonomous, in-person human–robot interaction. The sole experimental factor was the robot’s interaction policy. Participants were initially randomly assigned to condition; however, final condition allocation was partially constrained by participant attrition and scheduling, as described below.
Throughout this paper, references to “the robot” denote the autonomous interactive system comprising the Misty-II hardware platform and an integrated offboard software pipeline. Spoken-language understanding, dialogue management, task logic, and interaction policy execution were handled on an external edge device which interfaced with the robot via websocket interfaces. The Misty-II platform was responsible for audio capture, text to speech, and the execution of embodied behaviours including facial expressions, body movement, and LED signalling. Despite this distributed execution, all interaction decisions were generated autonomously by an AI driven centralized dialogue agent responsible for coordinating spoken-language understanding, task state, and verbal and nonverbal behaviour based on the interaction policy. More detail on the robot platform and software architecture is provided in Appendix A.
2.1 Interaction Policies
Across both tasks, interaction behaviour was governed by one of two interaction policies that differed in how the robot was intended to initiate, frame, and adapt its contributions during collaboration. Under both policies, the robot continuously monitored interaction timing and issued brief check-ins following extended periods of participant silence in order to preserve interaction continuity. In the neutral, reactive condition, these check-ins were minimal and task-focused, serving only to signal availability without providing guidance, encouragement, or additional framing (e.g., “I’m ready for your next question.”).
In contrast, under the responsive policy, the robot’s utterances were affectively framed and context-sensitive. In addition to answering questions, the robot proactively adapted its dialogue, assistance, and repair strategies based on inferred interaction state, such as hesitation, frustration, or apparent difficulty. This included acknowledging task difficulty, offering encouragement, and proposing collaborative reasoning (e.g., “I can tell you’re frustrated, don’t worry! we can reason through this together”), rather than waiting for an explicit request for help. Beyond differences in check-in style, the responsive robot also initiated guidance or clarification when interaction stagnated, whereas the neutral robot limited its contributions to basic task guidance and direct queries.
Both conditions operated fully autonomously without human intervention and otherwise used the same underlying task logic and sensing infrastructure.
2.2 Collaborative Task Design
The task structure was modelled after [25] and designed to examine collaboration under two distinct dependency conditions: enforced collaboration, in which successful task completion required the robot’s involvement, and optional collaboration, in which participants could choose whether and how to engage the robot. To operationalize these conditions in a naturalistic way, we embedded both tasks within a single immersive, narrative-driven puzzle scenario.
Participants were positioned as investigators searching for a missing laboratory robot named Atlas. The Misty-II robot served as a diegetic guide and collaborative partner throughout the session, framing each stage of the interaction as part of the unfolding investigation. This narrative structure provided continuity across tasks while allowing the dependency manipulation to vary between stages without disrupting immersion. The full session lasted approximately 25 minutes and consisted of five sequential stages, including two timed reasoning tasks.
Interactions took place in a shared physical workspace that included the Misty-II social robot and a participant-facing computer interface [19]. The interface displayed task materials, recorded participant responses, and supported progression through the game (see Figure 2). Importantly, it did not function as a control mechanism for the robot. Instead, the robot autonomously monitored task progression and participant inputs via the interface, adapting its dialogue and behaviour accordingly.
Figure 1: Experimental setup showing the autonomous robot and participant-facing task interface used during in-person sessions. Participants entered task responses and navigated between task stages using the interface, while the robot autonomously tracked task state and adapted its interaction based on participant input.
The interaction began with a brief greeting phase during which the robot introduced itself and engaged in rapport-building dialogue (e.g., “Hi, I’m Misty. What’s your name?”). This was followed by a mission briefing outlining the investigative scenario and establishing the collaborative goal: to determine who had taken Atlas and where it might now be located.
Participants then completed two core task stages. The first was a robot-dependent collaborative reasoning task in which Misty’s participation was required to solve the problem. The second was a more open-ended and complex problem-solving task in which robot assistance was available but not required. The interaction concluded with a wrap-up stage in which the robot provided closing feedback and formally ended the session.
Participants advanced between stages using the interface, either in response to the robot’s prompts or at their own discretion. All spoken dialogue and interaction events were managed autonomously by the robot and logged automatically for analysis.
2.2.1 Task 1: Robot-Dependent Collaborative Reasoning
In the first task, participants were asked to identify who had taken the missing robot from a 6 × 4 grid of 24 potential “suspects.” Within the narrative, Misty informed participants that she had access to internal security footage but was unable to directly reveal the suspect’s identity due to laboratory security protocols. Instead, she could answer only yes/no questions about observable features of the individual in the footage (e.g., “Was the suspect wearing a hat?”). This constraint established the robot as the sole source of task-relevant information while maintaining narrative plausibility.
The suspect grid was displayed on the interface and allowed participants to click on and grey out eliminated candidates, while all questions were posed verbally to the robot (see Figure 2).
Figure 2: Task 1 interface including the 6 × 4 grid of 24 candidates. Participants could track those eliminated by clicking on subjects which would grey them out. A box was provided to input their final answer and a button included to move to the next task.
The robot possessed the ground-truth information necessary to answer each question correctly. Successful task completion was therefore fully dependent on interaction with the robot, creating an enforced collaborative dynamic. Participants were required to coordinate a questioning strategy that progressively eliminated candidates based on shared features, narrowing the search space within a five-minute time limit. Efficient performance depended on selecting informative questions (i.e., features that divided the remaining candidates), tracking eliminations, and adapting subsequent questions based on prior answers.
This structure made the task sensitive to interaction quality. Inefficient questioning, repeated queries, or uncertainty about next steps could slow progress and increase cognitive load, whereas effective coordination supported rapid elimination and convergence on a solution. The constrained yes/no format ensured consistent informational bandwidth across participants and conditions, while still allowing meaningful variation in collaboration strategy. The robot’s behaviour under the responsive policy was designed to support this process by providing adaptive guidance, encouragement, and clarification when participants appeared to struggle or become disengaged, whereas the neutral robot provided only direct answers without additional support.
2.2.2 Task 2: Open-Ended Collaborative Problem Solving
The second task involved a more open-ended reasoning challenge designed to shift the dependency structure from enforced to optional collaboration. Participants were presented with multiple technical logs through a simulated terminal interface that could be used to infer the location of the missing robot (see Figure 3). The task was intentionally cryptic and difficult to solve within the allotted ten minutes without synthesizing information across multiple sources.
Figure 3: The task 2 interface presented multiple technical logs through a simulated terminal interface that could be used to determine the location of the missing robot.
The logs contained partial and indirect clues related to the robot’s activity, such as wireless connectivity records, sensor readings, and timestamped system events. Solving the task required participants to identify which logs were relevant, extract spatial or temporal cues, and integrate these signals to progressively narrow down plausible locations. As in Task 1, successful performance depended on managing uncertainty and iteratively refining hypotheses rather than on recognizing a single explicit cue.
Critically, unlike Task 1, the robot did not possess ground-truth information about the solution or direct access to the contents of the logs. This constraint was intentional: Misty’s assistance was limited to general reasoning support derived from its language model, such as explaining how to interpret log formats, suggesting strategies for cross-referencing timestamps, or prompting participants to reconcile inconsistencies across sources. Participants could complete the task independently or solicit assistance from the robot at their discretion [25], allowing collaboration to emerge voluntarily rather than being structurally enforced.
2.3 Study Protocol
Participants signed up for the study and completed a pre-session questionnaire before their in-person session via Qualtrics. The elapsed time between sign-up and the in-person session was at most week and at minimum immediately before the session. The pre-session questionnaire colleced basic demographics information and assessed baseline characteristics, including the Negative Attitudes Toward Robots Scale (NARS) and the short form of the Need for Cognition scale (NFC-s) [32, 7]. These measures were used to capture a baseline of individual differences that may moderate responses to robot interaction.
In-person sessions were conducted in a quiet, private room at Laurentian University between November and December 2025. Prior to each session, the robot’s interaction policy was configured to the assigned experimental condition.
Upon arrival, participants were greeted by the researcher, provided with a brief overview of the session, and given instructions for effective communication with the robot. Once participants indicated readiness, the researcher exited the room, leaving the participant and robot to complete the interaction without human presence or observation. Participants initiated the interaction by clicking a start button on the interface and were informed that they could terminate the session at any time without penalty.
Following task completion, participants completed a 21-item post-interaction questionnaire assessing trust on a laptop. Participants then engaged in a brief debrief with the researcher and were awarded a \$15 gift card. Total session duration averaged approximately 30 minutes.
2.4 Measures
A combination of self-report and objective measures was used to assess trust, engagement, and task performance.
2.4.1 Self-Report Measures
Participants completed a pre-session questionnaire assessing baseline characteristics, including the Negative Attitudes Toward Robots Scale (NARS) and the short form of the Need for Cognition scale (NFC-s) [7, 32]. These measures were used to capture a baseline of individual differences that may moderate responses to robot interaction.
Trust was assessed using two established self-report instruments commonly used in human–robot interaction research: the Trust Perception Scale–HRI (TPS-HRI) and the Trust in Industrial Human–Robot Collaboration scale (TI-HRC) [34, 9]. Both measures were adapted to reflect the specific dialogue-driven task context and interaction modality of the present study, while preserving the original constructs and response intent of each scale. 9 items were retained from the TI-HRC and 12 items from the TPS-HRI. Item wording was modified to reference the robot’s behaviour during the dialogue-driven collaborative tasks and response formats were adjusted to ensure interpretability for participants without prior robotics experience (see Appendix B for a full item list). Internal consistency of the adapted trust measures was evaluated within the present sample to assess measurement reliability under the modified wording and response formats. Given the pilot sample size, reliability estimates should be interpreted cautiously, but are reported to provide transparency regarding scale performance under fully autonomous interaction conditions.
All self-report scales demonstrated acceptable to strong internal consistency in the present sample. The Negative Attitudes Toward Robots scale (NARS) showed good reliability ($\alpha$ = .81), with subscale alphas ranging from .72 to .76. The Need for Cognition scale (NFC) demonstrated strong internal consistency ($\alpha$ = .87). Both post-interaction trust measures also exhibited high reliability. The Trust Perception Scale–HRI (TPS-HRI) showed strong internal consistency overall ($\alpha$ = .89), with subscale reliabilities ranging from .74 to .88. The Trust in Industrial Human–Robot Collaboration scale (TI-HRC) demonstrated very high internal consistency ($\alpha$ = .95).
Together, these instruments capture complementary dimensions of trust, including perceived reliability, task competence, and affective comfort. However, they differ in their conceptual emphasis: the TPS-HRI primarily operationalizes trust as a reflective judgement of system performance (i.e., “What percent of the time was the robot reliable”), whereas the TI-HRC scale emphasizes trust as an experienced, embodied response arising during interaction (i.e., “The way the robot moved made me feel uneasy”). Despite this complementarity, both measures rely on retrospective self-report and may be insensitive to moment-to-moment trust dynamics as collaboration unfolds. For this reason, questionnaire data were interpreted alongside behavioural and interaction-level measures.
2.4.2 Objective Measures
Table 3 outlines objective task metrics including the average number of dialogue turns, session duration, response time (robot dialogue latency), as well as the number of engaged versus frustrated responses the robot detected–all obtained from log data. Behavioural engagement metrics were derived from interaction logs and manually coded dialogue transcripts, including the proportion of turns characterized by communication breakdowns, human and robot reasoning, robot helpful and unhelpful guidance, robot and human affective engagement and other task-relevant human and robot contributions. See Appendix C for a full coding scheme.
2.5 Participants
A total of 29 participants were recruited from the Laurentian University community via word of mouth and the SONA recruitment system. Eligibility criteria required being 18 years or older, fluent in spoken and written English, and having normal or corrected-to-normal hearing and vision. Participants received a \$15 gift card as compensation for their time; some students additionally received partial credit for participating depending on their program of study. All procedures were approved by the Laurentian University Research Ethics Board (REB #6021966).
Although English fluency was an eligibility requirement, in-person observation during data collection revealed meaningful variability in participants’ functional spoken-language proficiency. The researcher therefore documented observed language proficiency and monitored interaction quality during each session in anticipation of potential speech-based system limitations.
Subsequent review of interaction transcripts and system logs indicated that a subset of sessions exhibited severe and sustained communication failure, characterized by fragmented or unintelligible automatic speech recognition (ASR) output and stalled dialogue. In these sessions, the robot was unable to extract sufficient linguistic content to maintain conversation, respond meaningfully to participant input, or support task progression. These interaction failures are described in detail in the Analytic Strategy and Results sections.
2.5.1 Randomization Check
Participants were initially assigned to the responsive or control condition by Qualtrics using an automated randomization procedure at the time of online sign-up. However, because several scheduled participants did not attend their in-person session, replacement participants were assigned to the next available session slot rather than re-randomized. As a result, the final condition assignment should be considered semi-random rather than strictly randomized.
To assess whether this deviation introduced systematic bias, we conducted randomization checks comparing participants across conditions on key baseline variables. Across analyses, participants in the responsive and control conditions were comparable with respect to demographic characteristics, prior experience with robots, and baseline attitudes toward robots, including gender, age, Negative Attitudes Toward Robots (NARS), and Need for Cognition scores (see Table 1; [7]). These patterns were consistent across both the eligible and full analytic samples, suggesting that the final group composition remained well balanced despite the partial breakdown of the intended randomization procedure.
Table 1: Baseline participant characteristics by interaction policy (full sample)
Characteristic N CONTROL N = 131 RESPONSIVE N = 161 p-value2
Gender 27 0.84
      Woman 6 / 13 (46%) 7 / 14 (50%)
      Man 7 / 13 (54%) 7 / 14 (50%)
Age Group 27 0.35
      18-24 5 / 13 (38%) 7 / 14 (50%)
      25-34 4 / 13 (31%) 2 / 14 (14%)
      34-44 1 / 13 (7.7%) 4 / 14 (29%)
      45+ 3 / 13 (23%) 1 / 14 (7.1%)
Program 25 >0.99
      Psychology 1 / 13 (7.7%) 1 / 12 (8.3%)
      Engineering 2 / 13 (15%) 1 / 12 (8.3%)
      Computer Science 7 / 13 (54%) 6 / 12 (50%)
      Earth Sciences 0 / 13 (0%) 1 / 12 (8.3%)
      Other 3 / 13 (23%) 3 / 12 (25%)
Experience with Robots 29 7 / 13 (54%) 4 / 16 (25%) 0.14
Native English Speaker 29 0.53
      Native English 5 / 13 (38%) 8 / 16 (50%)
      Non-Native English 8 / 13 (62%) 8 / 16 (50%)
NARS Overall 29 38 (8) 38 (7) 0.79
Need for Cognition 29 3.62 (0.78) 3.74 (0.74) 0.55
Dialogue Viability 29 0.63
      exclude 3 / 13 (23%) 2 / 16 (13%)
      include 10 / 13 (77%) 14 / 16 (88%)
1n / N (%); Mean (SD)
2Pearson’s Chi-squared test; Fisher’s exact test; Wilcoxon rank sum test
2.6 Analytic Strategy
Because the study relied fundamentally on spoken-language collaboration, analyses were structured to explicitly account for interaction viability. Sessions were classified as non-viable when severe communication breakdown prevented sustained dialogue between the participant and the robot, rendering the experimental manipulation inoperative.
Communication viability was operationalized using a dialogue-level metric derived from manual coding of system logs. For each session, the proportion of dialogue turns affected by speech-recognition failure (i.e., fragmented or unintelligible utterances) was computed. Sessions in which more than 60% of dialogue turns were affected were classified as non-viable, reflecting cases in which spoken-language interaction could not be meaningfully sustained. This criterion closely aligned with sessions independently flagged by the researcher during administration. It should be noted that because communication viability may itself be influenced by interaction policy, excluding non-viable sessions may have introduced selection bias and inflated estimated policy effects.
For this reason, analyses were conducted using three complementary approaches. Primary analyses were performed on an eligible sample excluding non-viable sessions, reflecting interactions in which the spoken-language protocol and experimental manipulation operated as intended. Full-sample analyses including all sessions were conducted as sensitivity checks. Finally, mechanism-focused analyses compared viable and non-viable sessions on interaction-process metrics (e.g., ASR failure rates, dialogue turn completion, task abandonment) to characterize how severe communication breakdown alters interaction dynamics. Trust measures from non-viable sessions were not interpreted as valid estimates of human–robot trust under functional interaction, as the robot was unable to sustain dialogue or collaborative behavior in these cases.
All analyses were conducted using R (version 4.5.1) within the Quarto framework. Data manipulation and visualization utilized the tidyverse suite of packages [40], with mixed-effects models fitted using the lme4 and lmerTest packages [3, 23] while Bayesian hierarchical models were fitted using the brms package [6]. Summary tables were generated using the gtsummary package [36]. All code used for data processing and analysis is available at: GitHub Repository
3 Results
Of the 29 completed sessions, 5 met the pre-specified criterion for non-viable interaction due to severe and persistent communication failure. These sessions were characterized by high rates of ASR failure, incomplete dialogue sequences, and skipped task stages.
Primary results are therefore reported for the eligible sample (n = 24), with full-sample and mechanism-focused analyses reported as sensitivity and exploratory analyses, respectively.
3.1 Primary Analysis: Eligible Sample
Simple descriptive comparisons of post-interaction trust measures indicated higher trust ratings in the responsive condition relative to the control condition across both trust scales (see Table 2). As indicated in Figure 4, average raw post-interaction scores on the Trust in Industrial Human–Robot Collaboration scale (TI-HRC) differed by approximately 26 points (Likert 1-5 converted to 0-100 scale for easier comparison across scales), while differences on the Trust Perception Scale–HRI (TPS-HRI) were approximately 15 points higher in the responsive condition compared to the control (Figure 4).
Importantly, overall task accuracy did not differ significantly between conditions (detailed task performance results are reported below), suggesting that observed differences in trust were not driven by differential task success but rather by the quality of the interaction process itself.
Table 2: Post-Interaction Trust Outcomes and Task Accuracy by Policy in the Eligible Sample
Characteristic CONTROL N = 101 RESPONSIVE N = 141 p-value2
Experienced Trust (TI-HRC) 39 (22) 67 (21) 0.004
Subscales
      Reliability 40 (24) 65 (18) 0.012
      Trust Perception 42 (23) 60 (22) 0.075
      Affective Trust 50 (31) 79 (22) 0.018
Perceived Trust (TPS-HRI) 59 (17) 77 (18) 0.022
Overall Task Accuracy 0.60 (0.21) 0.66 (0.23) 0.47
1Mean (SD)
2Wilcoxon rank sum test; Wilcoxon rank sum exact test
Figure 4: Distribution of Trust Perception by interaction policy. Points represent individual observations; violins depict score distributions. Red points indicate group means with 95% confidence intervals. Statistical comparisons are reported in the Results section.
Analysis of dialogue interaction patterns confirmed successful manipulation of the interaction policies (see Table 3). Manual coding of dialogue transcripts revealed substantial and significant differences in robot behaviour across conditions (see Appendix C for coding scheme). In the responsive condition, the robot employed encouragement in 36% of dialogue turns compared to 0% in the control condition (p < .001), expressed empathy or acknowledged participant affect in 13% of turns versus 0% in the control (p < .001), and used collaborative language (e.g., “we,” “let’s”) in 42% of turns compared to 5% in control (p < .001). Interestingly, while both conditions included proactive check-ins following periods of participant silence, the control robot engaged in such check-ins at a higher rate (21% vs. 13% of turns, p = .033), consistent with its reactive policy that limited proactive assistance to structured silence monitoring.
Critically, proportions of communication breakdown did not differ between conditions (25% vs. 22%, p = .70), indicating that interaction policy did not systematically impact technical speech recognition viability among eligible participants. Participants in the responsive condition also exhibited higher levels of AI-detected engagement during interaction, with an average of 3.50 engaged responses (SD = 1.95) compared to 2.00 (SD = 2.21) in the control condition. Consistent with these dialogue differences, interactions in the responsive condition were characterized by longer session durations and slower robot response times, reflecting the latency and time associated with executing additional dialogue and affective support behaviours. As such, responsiveness in the present study reflects a bundled interaction profile rather than a single isolated manipulation. Observed trust differences may therefore partially reflect increased interaction exposure or dialogue richness in addition to affect-adaptive behavior per se. Future work should more finely disentangle these components.
Table 3: Post-Interaction Objective Measure Averages by Policy in the Eligible Sample
Characteristic CONTROL N = 101 RESPONSIVE N = 141 p-value2
Objective Measure Averages
      Dialogue Turns 34 (9) 33 (5) 0.45
      Session Duration (min) 13.24 (3.06) 15.26 (2.12) 0.084
      Response Time (ms) 14.37 (3.76) 17.24 (2.52) <0.001
      Silent Periods 5.60 (1.96) 4.71 (2.05) 0.29
      Engaged Responses 2.00 (2.21) 3.50 (1.95) 0.040
      Frustrated Responses 0.60 (0.70) 0.93 (1.21) 0.68
% of Dialogue Turns Characterized by…
      Comm. Breakdowns 0.25 (0.17) 0.22 (0.16) 0.70
Human Affective Engagement 0.04 (0.05) 0.10 (0.08) 0.032
      Human Reasoning 0.26 (0.12) 0.35 (0.16) 0.21
      Robot Reasoning 0.12 (0.07) 0.37 (0.14) <0.001
      Robot Helpful Guidance 0.68 (0.08) 0.84 (0.08) <0.001
      Robot Unhelpful Contributions 0.08 (0.04) 0.02 (0.03) 0.003
      Robot Encouragement 0.00 (0.00) 0.36 (0.11) <0.001
      Robot Empathy Expression 0.00 (0.00) 0.13 (0.09) <0.001
      Robot Collaborative Language 0.05 (0.05) 0.42 (0.16) <0.001
      Robot Proactive Check-ins 0.21 (0.08) 0.13 (0.08) 0.033
1Mean (SD)
2Wilcoxon rank sum test; Wilcoxon rank sum exact test
Beyond descriptive comparisons, hierarchical models were fitted to evaluate the robustness of interaction policy effects while controlling for baseline covariates and accounting for measurement structure.
3.2 Hierarchical Models
Given the pilot nature of the study and the modest sample size, hierarchical modeling was used to estimate effect direction and associated uncertainty rather than to support definitive confirmatory claims. A Bayesian framework was adopted to allow partial pooling across trust items and sessions, improving parameter stability under limited data. Results are therefore interpreted in terms of effect magnitude and posterior uncertainty rather than binary significance thresholds.
To operationalize this approach, mixed-effects models were fitted using both frequentist and Bayesian estimation frameworks. All models included interaction policy (responsive vs. control) as the primary fixed effect, with baseline negative attitudes toward robots (NARS) and native English fluency included as covariates. Random intercepts for sessions and trust items accounted for the repeated-measures structure. Frequentist models provided hypothesis tests against null effects, whereas Bayesian models quantified posterior uncertainty and enabled evaluation across varying levels of communication viability.
Model building proceeded by comparing a baseline model containing interaction policy alone against models incorporating theoretically motivated covariates. Adding NARS scores significantly improved model fit ($\chi^{2}$ = 4.82, p = .028), whereas prior experience with robots did not. Native English fluency did not significantly improve model fit but was retained due to its relevance for spoken-language interaction viability.
Bayesian models used weakly informative, scale-appropriate priors. The intercept was assigned a Normal(50, 25) prior, reflecting the 0–100 trust score range and centering estimates near the midpoint of the scale. Fixed effects were assigned Normal(0, 10) priors to constrain effects to plausible magnitudes. Random-effect and residual standard deviations were assigned Exponential(1) priors to regularize variance estimates while retaining flexibility. All models demonstrated satisfactory convergence ($\hat{R} \leq$ 1.01; effective sample sizes > 1000).
3.2.1 Primary Analysis: Eligible Sample (n = 24)
Frequentist linear mixed-effects models revealed consistent positive effects of responsive interaction on both trust measures. For experienced trust (TI-HRC), participants who interacted with the responsive robot reported significantly higher post-interaction trust than those in the control condition ($\beta$ = 16.28, SE = 5.14, t = 3.17, p = .005). Higher baseline negative attitudes toward robots were associated with lower trust scores ($\beta$ = -7.43, SE = 2.81, p = .016), while native English fluency was not significantly associated with trust. Inclusion of random intercepts for individual trust items improved model fit, indicating meaningful item-level variability beyond session-level differences. For perceived trust (TPS-HRI), a comparable pattern emerged, with responsive interaction associated with higher trust scores ($\beta$ = 14.17, SE = 6.50, t = 2.00, p = .046). However, random intercepts for trust items did not improve model fit for this scale, likely reflecting differences in scale format and response interface: TPS-HRI was administered using a continuous slider input via touchpad, whereas TI-HRC employed discrete Likert-style response options. The slider-based interface may have reduced response precision and item-level variability, though meaningful between-condition differences remained detectable at the aggregate level.
Bayesian estimation converged on similar effect magnitudes with high posterior certainty. As illustrated in Figure 5, for TI-HRC, the responsive condition showed a posterior median effect of $\beta$ = 14.98 (95% credible interval [7.29, 22.22]), with near-unity probability of a positive effect. For TPS-HRI, the posterior median was $\beta$ = 12.76 (95% credible interval [2.96, 22.06]), with posterior probability exceeding 99% that the effect was positive. Baseline NARS showed credible negative associations with both outcomes, while native English fluency showed negative associations that were credible for TPS-HRI but uncertain for TI-HRC. Model fit was substantial for TPS-HRI (conditional $R^{2}$ = 0.64) and moderate for TI-HRC (conditional $R^{2}$ = 0.42), with fixed effects explaining 16% and 21% of variance, respectively. The smaller item-level variance for TI-HRC suggests greater coherence among affective trust items under functional interaction conditions.
Figure 5: Posterior distributions of fixed effects from the final Bayesian mixed-effects models (eligible sample: n=24) predicting post-interaction trust. Fixed effects include Responsive vs Control policy, Negative Attitudes Towards Robots (NARS) and Non-native English speaker versus English native speaker. Half-eye densities show posterior distributions; points indicate posterior medians; thick and thin intervals denote 80% and 95% credible intervals, respectively. The dashed vertical line marks a null effect, and positive values indicate higher trust scores on the respective scale (points on a 0–100 scale).
3.2.2 Sensitivity Analysis: Full Sample (n = 29)
To assess robustness, Bayesian models were refitted including all sessions regardless of communication viability. The responsive interaction effect remained positive for both trust measures but showed substantial attenuation compared to the eligible sample. For TPS-HRI, the posterior median effect was $\beta$ = 7.04 (95% credible interval [-1.83, 15.67]). Although uncertainty increased and the credible interval included zero, the posterior probability of a positive effect remained high (>94%). Model fit decreased relative to the eligible sample (conditional $R^{2}$ = 0.44), indicating increased unexplained variability when sessions with severe communication breakdown were included. For TI-HRC, attenuation was more pronounced. The posterior median effect decreased to $\beta$ = 7.17 (95% credible interval [-1.97, 16.70]), with reduced probability of a large effect. Model fit remained moderate (conditional $R^{2}$ = 0.60), but residual variance increased, consistent with the inclusion of interactions in which collaborative behaviour could not be sustained. These results indicate that experienced trust is particularly sensitive to interaction breakdown, and that trust ratings obtained under non-functional interaction conditions do not reflect graded variation in collaborative experience but rather reflect the collapse of the interaction itself.
3.2.3 Mechanism Analysis: Communication Breakdown (n=29)
To examine whether communication quality moderated the effect of interaction policy, Bayesian models were fitted in the full sample with the proportion of turns characterized by communication breakdown included as an interaction term. For TPS-HRI, the interaction between policy and communication breakdown was weak and centered near zero (posterior median $\beta$ = -1.14, 95% credible interval [-18.87, 16.28]), with only 55% probability of being negative and 50% probability of being significant. This broad, unstable posterior distribution suggests that evaluative trust judgments were relatively insensitive to graded variation in communication quality once a basic threshold of viability was established. In contrast, TI-HRC showed a more consistent pattern (see Figure 6). The interaction term showed a negative tendency (posterior median $\beta$ = -5.97, 95% credible interval [-23.01, 10.92]), with 76% probability of being negative, 71% probability of being significant, and 45% probability of being large. While responsive behaviour was associated with higher experienced trust under low levels of communication breakdown, this advantage diminished as communication failures accumulated. This differential sensitivity suggests that experienced trust depends critically on the robot’s sustained ability to engage in responsive interaction, whereas evaluative trust judgments may rely more on discrete moments of successful collaboration that can occur even within partially degraded interactions.
Figure 6: Predicted trust scores as a function of communication breakdown, stratified by interaction policy. Panels show experienced trust (left) and perceived trust (right). Lines represent posterior medians and shaded bands indicate 80% credible intervals. Communication breakdown is defined as the proportion of interaction turns flagged as breakdowns; predictions are shown at mean baseline negative attitudes toward robots.
3.3 Task performance
Overall task accuracy did not differ significantly between conditions (60% vs. 66%, p = .534), nor did performance on individual task components reach statistical significance (see Table 4). However, a notable pattern emerged when examining task structure. For the suspect identification task (Task 1), which required robot collaboration to complete accurately, participants in the responsive condition achieved more than double the accuracy of those in the control condition (64% vs. 30%, p = .106), representing a large effect that approached but did not reach conventional significance thresholds given the pilot sample size. In contrast, performance on the location identification task components (building, zone, and floor identification), where robot assistance was optional, showed no consistent directional advantage for either condition. These patterns suggest that responsive robot behaviour may particularly benefit collaborative performance on tasks requiring sustained interaction and mutual grounding, though larger samples are needed to establish statistical reliability. Critically, the absence of significant overall accuracy differences indicates that observed trust differences cannot be attributed simply to differential task success, but rather reflect distinct responses to the interaction process itself.
Table 4: Post-Interaction Task Performance by Policy in the Eligible Sample
Characteristic CONTROL N = 101 RESPONSIVE N = 141 p-value2
% Task Accuracy 0.60 (0.21) 0.66 (0.23) 0.47
Suspect ID Accuracy (robot dependent) 3 / 10 (30%) 9 / 14 (64%) 0.10
Building ID Accuracy 7 / 10 (70%) 11 / 14 (79%) 0.67
Zone ID Accuracy 5 / 10 (50%) 4 / 14 (29%) 0.40
Floor ID Accuracy 7 / 10 (70%) 13 / 14 (93%) 0.27
1Mean (SD); n / N (%)
2Wilcoxon rank sum test; Pearson’s Chi-squared test; Fisher’s exact test
3.4 Individual differences and correlational patterns
The following correlational analyses are exploratory in nature. A broad set of associations was examined across behavioural, interactional, and self-report measures, and p-values were not adjusted for multiple comparisons. These analyses are therefore intended to identify potentially meaningful patterns and generate hypotheses for future work rather than to provide confirmatory statistical tests.
Correlational analyses on the full sample (n=29) revealed patterns consistent with the interpretation that trust was shaped more strongly by interaction dynamics than by stable individual predispositions or objective task outcomes. A full correlation matrix is provided in Figure 7 to visualize the broader pattern of associations.
As expected, higher Need for Cognition (NFC) scores were negatively associated with baseline Negative Attitudes Towards Robots (NARS; r = -.48, p = .01), indicating that individuals who enjoy effortful thinking tend to hold more positive prior attitudes toward robots. However, neither NFC nor NARS showed significant associations with post-interaction trust outcomes (r = -.17 to -.33, p > .11), suggesting that the impact of responsive robot behaviour was relatively independent of participants’ baseline dispositions toward robots or cognitive engagement preferences. At the same time, higher NFC scores were strongly associated with greater engagement in reasoning dialogue during collaboration (r = .57, p = .003), indicating that cognitive engagement tendencies may have influenced interaction behaviour even if they did not directly determine trust evaluations (see Figure 7).
Consistent with the experimental manipulation, specific robot dialogue behaviours showed substantial correlations with trust outcomes. Robot empathy expression—quantified as the proportion of dialogue turns in which the robot acknowledged participant affect or expressed understanding—was strongly correlated with both experienced trust (TI-HRC; r = .53, p = .008) and perceived trust (TPS-HRI: r = .65, p = .001). Similarly, robot use of collaborative language (e.g., “we,” “let’s”) was positively associated with experienced trust (TI-HRC: r = .50, p = .012), as was robot encouragement (TI-HRC: r = .45, p = .027; TPS-HRI: r = .42, p = .040). These associations provide evidence that the specific affective and collaborative behaviours implemented in the responsive condition were linked to participants’ trust evaluations beyond simple condition assignment.
Figure 7: Correlation Matrix of Key Study Variables
Participant engagement during interaction, operationalized as the frequency of detected positive affective responses, was positively correlated with perceived trust (r = .52, p = .009) and showed a trending association with experienced trust (r = .30, p = .156). Engaged responses (both AI detected and affective dialogue by the human) were associated with longer interaction duration (r = .54, p = .007), greater use of collaborative language by the robot (r = .50, p = .013), and fewer communication breakdowns (r = -.49, p = .016). This pattern suggests that responsive robot behaviour may have fostered a reciprocal dynamic in which robot affective adaptation elicited participant engagement, particularly in those high in NFC, which in turn supported smoother interaction and higher trust.
Notably, objective task performance showed no significant association with either full trust measure (suspect accuracy: r = .18, -.20, p > .35; overall accuracy: r = -.05 to .19, p > .37), but did show a significant negative association with the TI-HRC Reliability subscale (r = .40, p < .05), as well as a negative association with detected frustration, silent periods and communication breakdowns. The dissociation between task outcomes and overall trust ratings suggests that participants’ trust judgments primarily reflected the quality and affective tone of the interaction process rather than instrumental task success. Task performance was not included as a covariate in primary models to avoid conditioning on a potential mediator of interaction policy effects.
4 Discussion
This pilot study examined how a responsive interaction policy influences trust during embodied, fully autonomous, spoken-language human–robot collaboration. Unlike much prior work in HRI trust research, all dialogue management, affect inference, task progression, and response generation were executed autonomously and in real time, without Wizard-of-Oz control or scripted repair. Participants were therefore exposed not only to adaptive interaction behaviour, but also to the latency, recognition errors, and coordination failures characteristic of deployed autonomous systems. The study was conducted as a pilot, with the primary aim of assessing feasibility and clarifying mechanisms under these realistic constraints rather than providing definitive confirmatory tests.
Across analytic approaches, the responsive interaction policy was associated with higher post-interaction trust compared to a neutral, reactive control policy when interaction remained viable. These differences emerged without reliable differences in overall task accuracy. Although the robot-dependent task showed a descriptive advantage for the responsive condition, this effect did not reach statistical significance in the present pilot sample. Taken together, this pattern suggests that trust judgments were shaped more strongly by the quality and tone of the interaction process than by instrumental task success alone. This is consistent with theoretical accounts that conceptualize trust as an emergent property of interaction dynamics rather than a static attribute of the system or a simple function of performance outcomes [24, 13].
Trust in human–robot interaction is widely recognized as multidimensional, encompassing perceptions of reliability and competence as well as affective comfort and perceived collaboration [24, 13]. The use of two validated trust measures in the present study allowed these components to be examined in parallel. Rather than treating trust as a static attribute of the system, the results suggest that it functioned as an evaluation of the collaborative experience as it unfolded over time. Notably, affective and experiential components of trust were more sensitive to interaction degradation than reliability-focused evaluations, reinforcing the view that trust in autonomous HRI is shaped by process-level interaction dynamics rather than performance outcomes alone.
These differences became particularly apparent under conditions of reduced communication viability. When speech recognition failures accumulated and conversational grounding could not be re-established, the trust advantage associated with responsiveness diminished. In these degraded sessions, the responsive policy continued generating proactive assistance and repair-oriented behaviour. However, when linguistic grounding could not be restored, these behaviours did not reliably recover collaboration and may instead have increased cognitive load relative to the neutral policy’s more restrained style. This pattern suggests that affect-responsive policies operate within a viability window: they enhance trust when grounding is intact but cannot compensate for sustained communicative breakdown.
A key aspect of this study is that these dynamics were observed under full autonomy. Because failures were not covertly repaired by a human experimenter, communication viability became an explicit feature of the interaction rather than a hidden nuisance variable. This visibility clarifies an important constraint: trust formation in spoken-language HRI presupposes sufficient linguistic grounding. When that grounding collapses, trust ratings may no longer reflect calibrated evaluations of competence or benevolence, but instead the breakdown of the collaborative process itself.
Taken together, the findings support a process-oriented account of trust in autonomous HRI. Trust appears to emerge from how a system manages uncertainty, repair, and coordination over time, rather than from task performance alone. This pilot study identifies both promising effects of affect-responsive interaction policies under viable conditions and clear constraints imposed by communication breakdown. These insights provide a concrete foundation for the design of subsequent studies aimed at stabilizing language grounding and more precisely isolating the mechanisms through which responsiveness shapes collaborative experience.
5 Limitations
Several limitations constrain interpretation of the present findings and point directly to areas for refinement.
The study was conducted as a pilot with a small sample size, limiting statistical power for detecting higher-order interactions involving task structure, communication quality, and individual differences. Although effect sizes were often substantial and broadly consistent across analytic approaches, uncertainty remains high. Replication with larger samples will be necessary to establish the stability of these effects and to more robustly test proposed mediation pathways.
Spoken-language interaction viability emerged as a central constraint, revealing an important direction for future research. Although English fluency was an eligibility requirement, substantial variability in functional spoken-language proficiency was observed during in-person sessions. In a subset of interactions, persistent speech recognition failure prevented the experimental manipulation from operating as intended, leading to exclusion from primary analyses on methodological grounds. Rather than reflecting a shortcoming of the experimental design, these cases highlight a fundamental challenge for autonomous spoken-language HRI: trust and collaboration presuppose a minimum level of linguistic grounding, and when that grounding fails, higher-level interaction constructs are no longer meaningfully instantiated.
The visibility of this constraint is itself a consequence of autonomy. In scripted or Wizard-of-Oz paradigms, language breakdown can be covertly repaired or masked by human intervention. In fully autonomous systems, communication viability becomes an explicit property of the interaction that must be detected, managed, and responded to by the robot. This points to a clear research direction focused on interaction policies that recognize emerging language mismatch and adapt accordingly, rather than assuming linguistic competence as a fixed prerequisite.
Natural language understanding was also constrained by the task-specific policy design used in the robot-dependent task. Fixed mappings between participant questions and predefined task features meant that semantically valid but unexpected phrasing, synonym use, or multi-attribute queries occasionally led to misinterpretation or incorrect responses (e.g., treating “orange hair” as distinct from the ground-truth feature “red hair”). These failures reflect limitations in prompt design and natural language understanding and robustness rather than participant reasoning and likely contributed to some interaction breakdowns.
Measurement-related factors may have introduced additional noise. One trust instrument relied on continuous slider-based responses administered via a laptop touchpad, whereas the other used discrete, clickable Likert-style responses. Touchpad-based slider interaction can be awkward and imprecise for some users, which may have attenuated effects on the continuous scale relative to the Likert-based measure. Future work should consider standardizing response modalities across trust measures to reduce measurement noise and improve sensitivity.
An additional limitation concerns the absence of a non-embodied comparison condition. The present study examined responsiveness within an embodied robotic platform, but did not include a functionally equivalent conversational agent presented through a simple chat interface. As a result, it remains unclear to what extent the observed trust effects are attributable to the interaction policy itself, the physical embodiment of the robot, or the interaction between embodiment and responsiveness. Future work should include a minimal conversational control condition in order to disentangle policy-level effects from embodiment-driven influences and to determine whether similar process-level trust dynamics emerge in purely virtual interaction contexts.
Finally, affect inference in the deployed system relied primarily on speech-based signals and conversational context. This design choice prioritized real-time stability and robustness under autonomous constraints but necessarily limited the richness of affect sensing. Incorporating facial expression or prosodic features could improve responsiveness, though such approaches introduce additional latency, orchestration complexity, and failure modes that were beyond the scope of this pilot study.
6 Conclusions and Future Work
This pilot study provides evidence that affect-responsive interaction policies can influence trust during fully autonomous, in-person human–robot collaboration. The observed effects emerged independently of overall task accuracy, reinforcing the central role of interaction quality and affective responsiveness in shaping collaborative experience.
At the same time, the findings delineate clear boundary conditions. Trust advantages were most pronounced under viable interaction conditions and attenuated when communication breakdown accumulated. When communicative grounding collapsed, differences between interaction policies diminished and trust ratings became less interpretable as calibrated evaluations. Communication viability thus appears to function as a prerequisite for trust formation in spoken-language HRI rather than merely as a moderating variable.
These results directly inform the next phase of work at both empirical and systems levels. Larger samples will enable more precise estimation of effects and formal testing of mediation pathways linking responsiveness, interaction fluency, affective engagement, and differentiated trust components. Designs that better stabilize language handling will allow responsiveness to be evaluated under more controlled grounding conditions.
Architecturally, future iterations will move toward more explicit state-based or graph-oriented dialogue management, allowing conditional stage transitions, structured repair strategies, and flexible tool integration. Adaptive language management mechanisms will be explored to detect emerging communication mismatch, such as repeated recognition failures or repair loops, and to adjust linguistic complexity, pacing, or response format accordingly. These refinements are intended to expand the range of conditions under which affect-responsive policies can operate effectively.
More broadly, this pilot study underscores the importance of evaluating trust in autonomous systems under conditions that expose, rather than conceal, system limitations. Understanding how trust is shaped, constrained, and potentially restored under real autonomy conditions is essential for the responsible deployment of social robots in collaborative environments.